How many circular permutations are there given the numbers on the clock?



1
Expert's answer
2022-03-18T10:27:46-0400

Solution


There are two possibilities


  1. If we consider the clockwise rotation
  2. If we consider the counterclockwise rotation


If we consider both the rotations different from one another, then the total number of permutations is given by "(n-1)!"


In this case "n=12"


Therefore, the total permutations will be


"(12-1)!=11!"


"=11\times10\times9\times8\times7\times6\times5\times4\times3\times2\times1"


"=39916800"




Now if we consider both the rotations are not different from one another, then the total number of permutations is given by "=\frac{(n-1)!}{2!}"


In this case "n=12"


Therefore, the total permutations will be


"=\frac{(12-1)!}{2!}"


"=\frac{(11)!}{2!}"


"=19958400"